OpenStudy (anonymous):

Solve. Express the answer in simplest form. -1/3 +w = 5 3/8

OpenStudy (xmissalycatx):

w= 137 / 24

OpenStudy (dr0zier99):

w=137/24

OpenStudy (dr0zier99):

what she said

OpenStudy (anonymous):

can you help me with a few more

OpenStudy (dr0zier99):

i can try

OpenStudy (anonymous):



OpenStudy (dr0zier99):

b=58/5

OpenStudy (anonymous):

thx

OpenStudy (dr0zier99):

yep any more